# A Survey of Uncertainty Estimation Methods on Large Language Models

### v1

@@ A Survey of Uncertainty Estimation Methods on Large Language Models
Zhiqiu Xia Rutgers University                        Jinxuan Xu Rutgers University                        Yuqian Zhang Rutgers University                        Hang Liu Rutgers University
@@ 1 Introduction
Large Language Models (LLMs) have emerged as state-of-the-art solutions for a wide range of problems, mainly due to their unparalleled ability to generate coherent and contextually appropriate responses to diverse user prompts Ouyang et al. (2022); Zhao et al. (2024). However, with the increasing adoption of LLMs, concerns have grown regarding their tendency to produce biased, hallucinated, non-factual, and misaligned outputs Zhang et al. (2023); Huang et al. (2024b). These issues are further exacerbated by the fact that such flawed responses often appear highly fluent and convincingly realistic, making them difficult to detect.
A promising approach to addressing the challenge of misleading yet plausible responses is uncertainty estimation, which assigns an uncertainty or confidence score to the model’s output. Figure 1 provides an overview of this process. First, the LLM generates an initial response based on the input. Next, a confidence score is computed for this response. The score is then evaluated against a predefined threshold to determine the final output. If the confidence score meets or exceeds the threshold, the initial response is accepted; otherwise, the model outputs "I do not know," thereby reducing the risk of providing incorrect but convincingly realistic information to users.
@@
There is an urgent need for a comprehensive survey on LLM uncertainty estimation. Below, we highlight three of them: (i) Although uncertainty estimation has been extensively studied in traditional deep neural networks (DNNs)—with Bayesian and ensemble methods being notable examples Gawlikowski et al. (2023))—these techniques are not easily transferable to LLMs, due to the large number of parameters in LLMs. (ii) LLMs significantly transform society, creating a strong demand for a thorough study of uncertainty estimation tailored to LLMs. A survey of recent advances in LLM uncertainty estimation would provide a solid foundation for future development in the field. (iii) While there are three existing surveys on LLM uncertainty estimation, each has notable limitations. Specifically, Huang et al. (2024a) dedicates a substantial portion of its content to traditional DNN uncertainty estimation rather than focusing on LLMs. Geng et al. (2024) shifts its attention to uncertainty calibration and the applications of LLM uncertainty estimation, rather than providing a deep exploration of the core techniques. Similarly, Shorinwa et al. (2024) devotes much of its content to benchmarks and applications while lacking a complete view of the uncertainty estimation methods on LLMs.
This manuscript focuses on studying the uncertainty estimation methods within the context of LLMs, introducing a new taxonomy from the perspective of LLMs. We center our scope around techniques applicable during the inference stage. We emphasize the methods that do not require additional data Ren et al. (2023); Kumar et al. (2023); Tonolini et al. (2024) or model modifications Huang et al. (2023a); Liu et al. (2024), ensuring the broad applicability of this survey. Moreover, this is, to the best of our knowledge, the first survey that conducts a thorough evaluation of representative uncertainty estimation approaches across various datasets and domains. Built on the insights from our evaluations, we postulate two interesting future directions for LLM uncertainty estimation.
@@ 3.1 Problem Definition and Overview
Methods (§ 3.5), xnode, l sep=6mm, [Implicit clustering: Duan et al. (2024) Nikitin et al. (2024), tnode] [Explicit clustering: Kuhn et al. (2023) Farquhar et al. (2024) , tnode] ] [Consistency-based
Methods (§ 3.4), xnode, l sep=6mm, [Similarity-based: Huang et al. (2023b) Manakul et al. (2023) Gao et al. (2024) Chen and Mueller (2024) Zhang et al. (2024) Harsha Tanneru et al. (2024) Lin et al. (2024), tnode] [Agreement-based: Cole et al. (2023) Lyu et al. (2024) Hou et al. (2024), tnode]] [Latent Information
Methods (§ 3.3), xnode, l sep=6mm, [Hidden states-based: Chen et al. (2024) , tnode] [Probability distribution-based: Manakul et al. (2023) Jiang et al. (2023) Ahdritz et al. (2024), tnode] [Predicted probability-based: Jiang et al. (2021) Manakul et al. (2023) Kadavath et al. (2022) Portillo Wightman et al. (2023) Ling et al. (2024) Malinin and Gales (2021) Bakman et al. (2024) , tnode ]] [Verbalizing
@@
Figure 3: Taxomony of uncertainty estimation methods on LLMs.
@@
How to estimate uncertainty and confidence? As shown in Figure 2, for each input $\bm{x}$, an LLM model has an underlying response distribution for it ( 1). For ease of illustration, we assume the distribution is a normal distribution $N(\mu,\sigma^{2})$. Uncertainty estimation is to estimate the underlying variance $\sigma^{2}$. For example, the sample variance of $M$ different responses $\bm{r_{1}},\cdots,\bm{r_{M}}$ ( 2) can be an estimator for the variance, which indicates the variations of responses ($U$ in Figure 2).
There generally are two types of confidence, i.e., overall confidence $C$ and the confidence $C_{i}$ associated with each response candidate $\bm{r_{i}}$. The overall confidence $C$ is complementary to $U$, i.e., the precision $1/\sigma^{2}$ of the distribution is a confidence $C$ to the input. The associated confidence is related to $\bm{x}$ and the tokens in a specific response $\bm{r_{i}}$. To provide the final response to answer the input $\bm{x}$ given sampled responses, some literature resort to majority voting to select the most-voted response $\bm{r}_{*}$ ( 3Wang et al. (2023), while others choose to generate one extra response $\bm{r_{M+1}}$ with low-temperature settings ( 4) as Farquhar et al. (2024).
Surveyed papers overview. Figure 3 categorizes all the uncertainty estimation papers for LLM into four classes: verbalizing methods, latent information methods, consistency-based methods, and semantic clustering methods. We review each through Sections 3.23.5.
@@ 3.2 Verbalizing Methods
Lin et al. (2022a) pioneer this cohort of efforts. As the capabilities of LLMs continue to develop, they can provide reasonable confidence under proper guidance, even without fine-tuning. Subsequently, Tian et al. (2023) proposes three verbalizing variants: (i) Generate multiple response candidates with confidence scores and select the highest-rated one as the final response, (ii) derive the response and confidence through two rounds of prompt-and-answer interactions, and (iii) use words instead of numerical values to indicate the confidence. Recently, Harsha Tanneru et al. (2024) introduces two methods inspired by Chain-of-Thought (CoT) prompting. The first method requests the LLM to assign an importance score to each word in the input, while the second one prompts the LLM to provide confidence for each reasoning step in the response. Finally, LLM will offer a final confidence score for the overall response. Beyond that, Xiong et al. (2024) presents a systematic framework for verbalizing methods with three parts: prompting, sampling, and aggregation. It employs specific confidence-eliciting prompts and generates diverse response samples containing confidence scores. After that, the final confidence score is derived through inter-sample agreement or response ranking information.
@@ 3.3 Latent Information Methods
The analysis of token probabilities can be extended beyond a single response for more robust uncertainty estimation. Portillo Wightman et al. (2023) proposes to average the predicted probabilities across multiple responses. Ling et al. (2024) picks the key token from the responses and aggregates them into a distribution, and the uncertainty is from the entropy of the distribution. Kadavath et al. (2022) considers all the tokens in the responses, calculates the probability for each response using token probabilities, and measures uncertainty through the entropy of the response distribution, called predictive entropy. However, varying response lengths can introduce undesirable noise to the estimation. To address this limitation, Malinin and Gales (2021) proposes the length normalized entropy, incorporating the response length based on predictive entropy. Furthermore, Bakman et al. (2024) proposes to replace the length normalization by assigning a weight to each token with a BERT model to consider both the sequence length and semantic contribution of tokens.
While the methods above only require access to the probability value of the response tokens, the following papers would require access to the complete probability distributions: Manakul et al. (2023) computes the entropy of the probability distribution for each generated token, using either the mean or maximum entropy as the uncertainty. For multiple-choice questions, Jiang et al. (2023) presents a specialized methodology. It computes probability distributions over potential options for each response sample and aggregates these distributions to form an ensemble probability distribution for uncertainty estimation. Ahdritz et al. (2024) introduces a heuristic two-stage method. Initially, the LLM is prompted to generate multiple next-token candidates. Subsequently, through a "repeated prompt" mechanism, the model produces the next token. The final uncertainty score is then computed from the probability distribution of these next tokens.
Beyond the methods using the probability distributions of tokens in the response, some researchers utilize the hidden states of LLMs. Chen et al. (2024) proposes to use the embeddings in the middle layer of LLMs to construct a covariance matrix for responses, which captures the correlation relationships among them. By manipulating the eigenvalues of the covariance matrix, the degree of divergence among responses can be estimated and considered an uncertainty measure.
@@ 3.4 Consistency-based Methods
While the methods primarily focus on analyzing answer agreement patterns to estimate uncertainty, more methods emphasize evaluating the similarities among responses ( 3). For domain-specific tasks, targeted metrics like BLEU Papineni et al. (2002) and CodeBLEU Ren et al. (2020) have been successfully applied to machine translation and code generation tasks, respectively Huang et al. (2023b). In general question-answering scenarios, token-level similarity metrics such as BERTScore Zhang et al. (2020) and RougeL Lin (2004) have been widely adopted Huang et al. (2023b); Manakul et al. (2023); Gao et al. (2024). Moving beyond token-level comparisons, more sophisticated approaches that capture semantic relationships have emerged, including SentenceBERT and NLI-based methods Gao et al. (2024); Chen and Mueller (2024); Zhang et al. (2024). SentenceBERT computes the cosine similarity between two sentences using embeddings generated by the Sentence Transformer model. The NLI-based method leverages natural language inference (NLI) classifiers to categorize sentence relationships as entailment, neutral, or contradiction, regarding the probability the NLI classifier assigns to the “entailment” class as the similarity score. Moreover, Harsha Tanneru et al. (2024) proposes token importance uncertainty and CoT uncertainty. The former quantifies uncertainty through token agreement and token rank metrics, while the latter evaluates inter-step relationships using NLI classification techniques.
@@
While most methods estimate confidence by simply averaging similarities among responses in step ( 4), Lin et al. (2024) proposes a new similarity-based method for calculating confidence inspired by spectral clustering. It treats generated responses as nodes and obtains the degree matrix and the graph Laplacian matrix. Correspondingly, this method defines several uncertainty and confidence measures from the matrices.
@@ 3.5 Semantic Clustering Methods
Figure 7 depicts the workflow of semantic clustering methods, which leverages both the latent information and the semantic relationships among responses to offer a more comprehensive estimation of the uncertainty. The first two steps are similar to the consistency-based methods, where the LLM generates responses to the original input and its paraphrased versions ( 1- 2). Next, instead of calculating the similarities, the sampled outputs are partitioned into clusters with a new probability for each cluster ( 3). Finally, the probability distribution over these clusters calculates a confidence score ( 4). The motivation behind semantic clustering methods is that consistency-based methods regard two responses as consistent only if they have identical words, which is too strict. Therefore, semantic clustering of the latent information is proposed to deal with the limitations. We refer the readers to Section A.4 for the formula of different semantic clustering methods.
Kuhn et al. (2023) introduces semantic entropy for uncertainty estimation. The method comprises three phases: generation ( 1- 2), clustering ( 3), and entropy estimation ( 4). In step 3, a bi-directional entailment algorithm is employed to cluster semantically equivalent responses. It assesses the entailment relationship between each pair of responses, considering them to express the same meaning if they mutually entail each other. The entailment relationship can be determined with the help of an NLI classifier or by simply requesting a general-purpose LLM. The uncertainty is the entropy calculated from the cluster probabilities in step 4. In case there is no access to the token probability, Farquhar et al. (2024) introduces discrete semantic entropy, which leverages the response frequency to derive the aggregated probabilities.
While the above method clusters the responses explicitly, some methods propose implicit clustering. Instead of utilizing the bi-directional entailment algorithm, Duan et al. (2024) introduces sentence relevance scores between each response pair, which is more effective over long sentences than the bi-directional entail algorithm. Nikitin et al. (2024) further considers the distances between the clusters. The method encodes similarities among responses via positive semidefinite unit trace kernels. It offers a more fine-grained uncertainty measure using the von Neumann entropy of these kernels.

### v2

@@ A Survey of Uncertainty Estimation Methods on Large Language Models
Zhiqiu Xia, Jinxuan Xu, Yuqian Zhang, Hang Liu Rutgers, The State University of New Jersey {zhiqiu.xia, jinxuan.xu, yqz.zhang, hang.liu}@rutgers.edu
@@ 1 Introduction
Large Language Models (LLMs) have emerged as state-of-the-art solutions for a wide range of problems, mainly due to their unparalleled ability to generate coherent and contextually appropriate responses to diverse user prompts Ouyang et al. (2022); Zhao et al. (2024). However, with the increasing adoption of LLMs, concerns have grown regarding their tendency to produce biased, hallucinated, non-factual, and misaligned outputs Zhang et al. (2023b); Huang et al. (2024b). These issues are further exacerbated by the fact that such flawed responses often appear highly fluent and convincingly realistic, making them difficult to detect. A promising approach to addressing the challenge of misleading yet plausible responses is uncertainty estimation, which assigns an uncertainty or confidence score to the model’s output. Figure 1 provides an overview of this process. First, the LLM generates an initial response based on the input. Next, a confidence score is computed for this response. The score is then evaluated against a predefined threshold to determine the final output. If the confidence score meets or exceeds the threshold, the initial response is accepted; otherwise, the model outputs "I do not know," thereby reducing the risk of providing incorrect but convincingly realistic information to users.
@@
There is an urgent need for a comprehensive survey on LLM uncertainty estimation. Below, we highlight three of them: (i) Although uncertainty estimation has been extensively studied in traditional deep neural networks (DNNs)—with Bayesian and ensemble methods being notable examples Gawlikowski et al. (2023))—these techniques are not easily transferable to LLMs, due to the large number of parameters in LLMs. (ii) LLMs significantly transform society, creating a strong demand for a thorough study of uncertainty estimation tailored to LLMs. While numerous uncertainty estimation methods exist, recent benchmark studies Fadeeva et al. (2023); Vashurin et al. (2025) have focused on their empirical evaluation rather than in-depth methodological discussions. Consequently, a survey of recent advances in LLM uncertainty estimation that synthesizes recent progress is crucial, offering a solid foundation for future development in the field. (iii) Besides surveys focusing on natural language processing Fomicheva et al. (2020); Baan et al. (2023), there are three existing surveys concentrating on LLM uncertainty estimation. However, each has notable limitations. Specifically, Huang et al. (2024a) dedicates a substantial portion of its content to traditional DNN uncertainty estimation rather than focusing on LLMs. Geng et al. (2024) shifts its attention to uncertainty calibration and the applications of LLM uncertainty estimation, rather than providing a deep exploration of the core techniques. Similarly, Shorinwa et al. (2024) devotes much of its content to benchmarks and applications while lacking a complete view of the uncertainty estimation methods on LLMs.
This work focuses on studying the uncertainty estimation methods within the context of LLMs, introducing a new taxonomy from the perspective of LLMs. We center our scope around techniques applicable during the inference stage. We focus on the uncertainty estimation methods, excluding confidence calibration methods Zhou et al. (2023); Detommaso et al. (2024) from our scope. Besides, we emphasize the methods that do not require additional data Ren et al. (2023); Kumar et al. (2023); Tonolini et al. (2024) or model modifications Huang et al. (2023a); Liu et al. (2024), ensuring the broad applicability of this survey. Furthermore, this survey conducts a thorough evaluation of representative uncertainty estimation approaches across various datasets and domains. Built on the insights from our evaluations, we postulate two interesting future directions for LLM uncertainty estimation.
@@ 3.1 Problem Definition and Overview
Methods (§ 3.5), xnode, l sep=6mm, [Implicit clustering: Duan et al. (2024) Fadeeva et al. (2024) Lin et al. (2024) Nikitin et al. (2024), tnode] [Explicit clustering: Kuhn et al. (2023) Farquhar et al. (2024) , tnode] ] [Consistency-based
Methods (§ 3.4), xnode, l sep=6mm, [Similarity-based: Huang et al. (2023b) Manakul et al. (2023) Gao et al. (2024) Wang and Holmes (2024) Chen and Mueller (2024) Zhang et al. (2024) Harsha Tanneru et al. (2024) Lin et al. (2024), tnode] [Agreement-based: Cole et al. (2023) Lyu et al. (2024) Hou et al. (2024), tnode]] [Latent Information
Methods (§ 3.3), xnode, l sep=6mm, [Hidden states-based: Chen et al. (2024a) Sriramanan et al. (2024) , tnode] [Probability distribution-based: Manakul et al. (2023) Zhang et al. (2023a) Jiang et al. (2023) Ahdritz et al. (2024), tnode] [Predicted probability-based: Jiang et al. (2021) Manakul et al. (2023) Kadavath et al. (2022) Portillo Wightman et al. (2023) Ling et al. (2024) Malinin and Gales (2021) Bakman et al. (2024) , tnode ]] [Verbalizing
@@
Figure 3: Taxonomy of uncertainty estimation methods on LLMs.
@@
How to estimate uncertainty and confidence? Following the conceptualization by Lin et al. (2024), we illustrate the process as shown in Figure 2. For each input $\bm{x}$, an LLM model has an underlying response distribution for it ( 1). For ease of illustration, we assume the distribution is a normal distribution $N(\mu,\sigma^{2})$. Uncertainty estimation is to estimate the underlying variance $\sigma^{2}$. For example, the sample variance of $M$ different responses $\bm{r_{1}},\cdots,\bm{r_{M}}$ ( 2) can be an estimator for the variance, which indicates the variations of responses ($U$ in Figure 2).
There are generally two types of confidence, i.e., overall confidence $C$ and the confidence $C_{i}$ associated with each response candidate $\bm{r_{i}}$. The overall confidence $C$ is complementary to $U$, i.e., the precision $1/\sigma^{2}$ of the distribution is a confidence $C$ to the input. The associated confidence is related to $\bm{x}$ and the tokens in a specific response $\bm{r_{i}}$. To provide the final response to answer the input $\bm{x}$ given sampled responses, some literature resort to majority voting to select the most-voted response $\bm{r}_{*}$ ( 3Wang et al. (2023), while others choose to generate one extra response $\bm{r_{M+1}}$ with low-temperature settings ( 4Farquhar et al. (2024).
Survey papers overview. Figure 3 categorizes all the uncertainty estimation papers for LLM into four classes: verbalizing methods, latent information methods, consistency-based methods, and semantic clustering methods. We review each of them through Sections 3.23.5.
@@ 3.2 Verbalizing Methods
Lin et al. (2022a) pioneers this cohort of efforts. As the capabilities of LLMs continue to develop, they can provide reasonable confidence under proper guidance, even without fine-tuning. Subsequently, Tian et al. (2023) proposes three verbalizing variants: (i) Generate multiple response candidates with confidence scores and select the highest-rated one as the final response, (ii) derive the response and confidence through two rounds of prompt-and-answer interactions, and (iii) use words instead of numerical values to indicate the confidence. Recently, Harsha Tanneru et al. (2024) introduces two methods inspired by Chain-of-Thought (CoT) prompting. The first method requests the LLM to assign an importance score to each word in the input, while the second one prompts the LLM to provide confidence for each reasoning step in the response. Finally, LLM will offer a final confidence score for the overall response. Beyond that, Xiong et al. (2024) presents a systematic framework for verbalizing methods with three parts: prompting, sampling, and aggregation. It employs specific confidence-eliciting prompts and generates diverse response samples containing confidence scores. After that, the final confidence score is derived through inter-sample agreement or response ranking information.
@@ 3.3 Latent Information Methods
The analysis of token probabilities can be extended beyond a single response for more robust uncertainty estimation. Portillo Wightman et al. (2023) proposes to average the predicted probabilities across multiple responses. Ling et al. (2024) picks the key token from the responses and aggregates them into a distribution, and the uncertainty is from the entropy of the distribution. Kadavath et al. (2022) considers all the tokens in the responses, calculates the probability for each response using token probabilities, and measures uncertainty through the entropy of the response distribution, called predictive entropy. However, varying response lengths can introduce undesirable noise to the estimation. To address this limitation, Malinin and Gales (2021) proposes the length-normalized entropy, incorporating the response length based on predictive entropy. Furthermore, Bakman et al. (2024) proposes to replace the length normalization by assigning a weight to each token with a BERT model to consider both the sequence length and the semantic contribution of tokens.
While the methods above only require access to the probability value of the response tokens, the following papers would require access to the complete probability distributions: Manakul et al. (2023) computes the entropy of the probability distribution for each generated token, using either the mean or maximum entropy as the uncertainty. Zhang et al. (2023a) proposes an uncertainty metric that combines the sum of negative log probabilities and entropy, while also considering token importance, preceding context, and token properties such as entity type and token frequency. For multiple-choice questions, Jiang et al. (2023) presents a specialized methodology. It computes probability distributions over potential options for each response sample and aggregates these distributions to form an ensemble probability distribution for uncertainty estimation. Ahdritz et al. (2024) introduces a heuristic two-stage method. Initially, the LLM is prompted to generate multiple next-token candidates. Subsequently, through a "repeated prompt" mechanism, the model produces the next token. The final uncertainty score is then computed from the probability distribution of these next tokens.
Beyond the methods using the probability distributions of tokens in the response, some researchers utilize the hidden states of LLMs. Chen et al. (2024a) proposes to use the embeddings in the middle layer of LLMs to construct a covariance matrix for responses, which captures the correlation relationships among them. By manipulating the eigenvalues of the covariance matrix, the degree of divergence among responses can be estimated and considered an uncertainty measure. Different from this approach, Sriramanan et al. (2024) constructs a covariance matrix within a single response, where each row corresponds to a token in the response. The log-determinant of this covariance matrix is then calculated as an uncertainty metric. Additionally, it proposes another method based on the internal components of the attention mechanism. Specifically, it calculates the sum of the log-determinants of each self-attention head’s kernel similarity map, which serves as an alternative measure of uncertainty.
@@ 3.4 Consistency-based Methods
While the methods primarily focus on analyzing answer agreement patterns to estimate uncertainty, more methods emphasize evaluating the similarities among responses ( 3). For domain-specific tasks, targeted metrics like BLEU Papineni et al. (2002) and CodeBLEU Ren et al. (2020) have been successfully applied to machine translation and code generation tasks, respectively Huang et al. (2023b). In general question-answering scenarios, token-level similarity metrics such as BERTScore Zhang et al. (2020) and RougeL Lin (2004) have been widely adopted Huang et al. (2023b); Manakul et al. (2023); Gao et al. (2024). Moving beyond token-level comparisons, more sophisticated approaches that capture semantic relationships have emerged, including SentenceBERT and NLI-based methods Wang and Holmes (2024); Gao et al. (2024); Chen and Mueller (2024); Zhang et al. (2024). SentenceBERT computes the cosine similarity between two sentences using embeddings generated by the Sentence Transformer model. The NLI-based method leverages natural language inference (NLI) classifiers to categorize sentence relationships as entailment, neutral, or contradiction, regarding the probability the NLI classifier assigns to the “entailment” class as the similarity score. Moreover, Harsha Tanneru et al. (2024) proposes token importance uncertainty and CoT uncertainty. The former quantifies uncertainty through token agreement and token rank metrics, while the latter evaluates inter-step relationships using NLI classification techniques.
@@
While most methods estimate confidence by averaging similarities among responses in step ( 4), Lin et al. (2024) proposes using the number of semantic sets as a measure of uncertainty. These semantic sets are defined as “semantic equivalent” subsets, which are grouped from all responses using an NLI classifier.
Notably, consistency-based methods are computationally expensive due to the need for multiple inferences. While some approaches rely on auxiliary modules, such as an NLI classifier, their computational cost is considerably lower than that of performing an LLM inference. As a result, the computational overhead is primarily dominated by the number of inferences required.
@@ 3.5 Semantic Clustering Methods
Figure 7 depicts the workflow of semantic clustering methods, which leverages both the latent information and the semantic relationships among responses to offer a more comprehensive estimation of the uncertainty. The first two steps are similar to the consistency-based methods, where the LLM generates responses to the original input and its paraphrased versions ( 1- 2). Next, instead of calculating the similarities, the sampled outputs are partitioned into clusters with a new probability for each cluster ( 3). Finally, the probability distribution over these clusters calculates a confidence score ( 4). The motivation for semantic clustering methods stems from the limitations of consistency-based approaches, which can only deliver a perfect uncertainty score when the responses use identical wording. However, in reality, responses may convey the same meaning through varied expressions. Therefore, semantic clustering of the responses is proposed to deal with the limitations. We refer the readers to Section A.4 for the formula of different semantic clustering methods.
Kuhn et al. (2023) introduces semantic entropy for uncertainty estimation. The method comprises three phases: generation ( 1- 2), clustering ( 3), and entropy estimation ( 4). In step 3, a bi-directional entailment algorithm is employed to cluster semantically equivalent responses. It assesses the entailment relationship between each pair of responses, considering them to express the same meaning if they mutually entail each other. The entailment relationship can be determined with the help of an NLI classifier or by simply requesting a general-purpose LLM. The uncertainty is the entropy calculated from the cluster probabilities in step 4. In case there is no access to the token probability, Farquhar et al. (2024) introduces discrete semantic entropy, extending the work of Kuhn et al. (2023), which leverages the frequency of each cluster to calculate an entropy as the uncertainty.
While the above method clusters the responses explicitly, some methods propose implicit clustering. Duan et al. (2024) introduces sentence relevance scores between each response pair, which is more effective over long sentences than the bi-directional entail algorithm in the work of Kuhn et al. (2023). Fadeeva et al. (2024) proposes to further alleviate the impact of claim-type uncertainty by grouping words into several claim types. Besides, it only requires one-time inference by selecting the top-k choices of words using the latent information to generate different responses.
In contrast to these latent-information-dependent techniques, some semantic clustering methods operate without requiring access to such information. Lin et al. (2024) proposes to treat generated responses as nodes and obtain the degree matrix and the graph Laplacian matrix from the pairwise similarities of the responses. Correspondingly, this method defines several uncertainty and confidence measures from the matrices. Nikitin et al. (2024) further considers the distances between the clusters. The method encodes similarities among responses via positive semidefinite unit trace kernels. It offers a more fine-grained uncertainty measure using the von Neumann entropy of these kernels.